OpenStudy (anonymous):

2. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population.

OpenStudy (jack1):

"Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population."... are we missing part of the question, coz this is the first I'm reading about Iris or an Invasive snail population...?

OpenStudy (jack1):

does the function g(x) = 200(1.03)2x look like this: \[\huge g(x) = 200 \times (1.03)^(2x)\] ?

OpenStudy (jack1):

damn latex, standby

OpenStudy (jack1):

hmm... my equation function button isn't working either...

OpenStudy (anonymous):

I already got it, but thx.

OpenStudy (jack1):

sweet, nvmd thn

OpenStudy (anonymous):

aha, can you help me w another one tho?

OpenStudy (jack1):

sure, shoot

OpenStudy (anonymous):

3. Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing. &g(x) = 200(1.03)2x is the equation.

OpenStudy (jack1):

... u mean f(x) = 200(1.03)^2x... yeah (not g(x)) ...?

OpenStudy (anonymous):

yeah suree, lol

OpenStudy (jack1):

i would start at 200, and expand out to 1000 for the range, as you can then prove the time (x) for the population of the invasive species population to increase by 5 times...

OpenStudy (jack1):

domain would pretty much be all real numbers, as this equation would never be negative for the value of f(x)

OpenStudy (jack1):

but if you want to narrow it down, work out the x and y values at x = 0 (hint g(x) = 200, so there's your y-intercept) and at g(x) = 1000, (x=???). As this is an exponential growth function, and the base is greater than 1.0 (in this case = 1.03), it's a positive growth function, so its increasing as x increases and it's also getting steeper as x increases...

OpenStudy (jack1):

zat all make sense maddie? @madisonxo

OpenStudy (anonymous):

yes, thank you!

OpenStudy (jack1):

sweet as, hi five ;D